OpenStudy (anonymous):

Simplify? 4{-3+5[6-3(5+7)]}

OpenStudy (anonymous):

Ooo, fun. Small brackets first. =4{-3+5[6-3(12)]} =4{-3+5[6-36]} =4{-3+5[-30]} =4{-3-150} =4{-153} =-612 I think that's about right.

OpenStudy (anonymous):

Thank you miss